OpenStudy (anonymous):

A hockey puck slides off the edge of a table with an initial velocity of 20.0 m/s. The height of the table above the ground is 2.0m. What is the magnitude of the velocity of the puck just before it touches the ground? I know the answer is 21m/s but can you show me how?

OpenStudy (btaylor):

Kinematics shows us that the horizontal and vertical components of the motion are separable. The horizontal velocity (in this case, 20.0 m/s) is constant. The vertical velocity changes because of gravity's acceleration. The vertical velocity can be found with this equation (only because the initial vertical velocity is zero):\[v_f = \sqrt{2gh} = \sqrt{2 \times 9.8 \times 2} \approx 6.2\]Now that we know the vertical and horizontal components, we can find the resultant velocity:  Using Pythagorean's theorem: \[v_f = \sqrt{v_x^2 + v_y^2} = \sqrt{20^2 + 6.2^2} \approx 20.94\] So your answer is 20.94 m/s, which does indeed round to 21 m/s.

OpenStudy (anonymous):

You are awesome! Thanks!